Solve the equations. check the solutions.
sqrt(4x+1+3)=0

Found 2 solutions by oberobic, edjones:
Answer by oberobic(2304)

Squaring both sides, we have 4x + 1 + 3 = 0. Simplifying, 4x = -4, which means x = -1.

Answer by edjones(7569)

sqrt(4x+1+3)=0
4x+1+3=0 square both sides.
4x=-4 subtract 4 from each side.
x=-1 divide 4 into each side.
Check:
sqrt(-4+1+3)=0
sqrt(0)=0 true